a car begin to move from rest for the first half of path it moves with constant acceleration. then it moves with the attained velocity of 18m/s for the rest path.What is its average speed for the entire path ?

Grade:11

1 Answers

Niraj Bishnoi
6 years ago
Is the answer ~15.36m/s or 10.26m/s?
If yes then here’s how I did-
=>v2=v02+as                         (2xs/2=s)
=>v2=as …................................(A)
Now v becomes vfor the second case,
=>182=v2+as 
=>324=2as
=>as=162
=>v2=162
v=12.72
Now average speed, v (conjugate)= u+v/2=30.72/2=15.36
I don’t remember very well, but the average speed could also be=u+v+vf/3=30.72/3=10.26m/s